1-find the common ratio and the seventh term of the following sequence :2÷9,2÷3,6,18.
2- the first term of a geometric series is 1 and the fourth term is 27.find the sum of the first terms.
1
Expert's answer
2019-11-27T10:04:09-0500

1) The common ratio is

"q=\frac{18}{6}=3"

"b_7=b_1q^{7-1}=\frac{2}{9}3^{6}=162"

2)


"b_4=1q^{4-1}=q^3=27"

"q=3"

The sum of the first n terms:


"S_n=1\frac{1-3^n}{1-3}=\frac{3^n-1}{2}"